OpenStudy (anonymous):

What is the slope of the line that passes through (2,1) and (7,-2)? A: -3/5 B: -1/3 C: 1/5 D: 5/3

OpenStudy (alexandervonhumboldt2):

use slope formula (y1-y2)/(x1-x2)=(1+2)/(2-7)=3/(-5)=-3/5

OpenStudy (alexandervonhumboldt2):

slope formula is \[\frac{ y_1-y_2 }{ x_1-x_2 }\] or \[\frac{ y_2-y_1 }{ x_2-x_1 }\] you can use both formula's. there is no difference which to use.